OpenStudy (anonymous):

5.Consider the functions f → g∶ Z × Z → Z × Z defined as f (m,n) = (3m − 4n,2m + n) and g(m,n) = (5m + n,m). Find the formulas for g o f and f o g.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I'll do \(g\circ f\): Since \(f(m,n) = (3m-4n,2m+n)\) and \(g(m,n) = (5m+n,m)\), when we look at \(g(f(m,n))\), we plug in the first coordinate of \(f(m,n)\) wherever we had an \(m\) in \(g(m,n)\), and plug in the second coordinate of \(f(m,n)\) wherever we had an \(n\) in \(g(m,n)\). Thus, it follows that \[\begin{aligned} g(f(m,n)) &= g(3m-4n,2m+n)\\ & = (5(3m-4n)+(2m+n), (3m-4n))\\ &= (15m-20n+2m+n,3m-4n)\\ &= (17m-19n,3m-4n)\end{aligned}\] Do you think that you can do \(f\circ g\) following a similar process? I hope this made sense! :-)

OpenStudy (anonymous):

thank you.. i will try first for fog

OpenStudy (anonymous):

@ChristopherToni can u check for me.. f(g(m,n)) = f(5m + n,m) = (3(5m+n) + (2m + n), (5m + n)) = (15m + 3n + 2m +n , 5m + n ) = (17m + 4n, 5m + n) is tht true ?

OpenStudy (anonymous):

Unfortunately, you didn't evaluate \(f(5+n,m)\) correctly. :/ When doing this, think of it as letting \(s=5m+n\) and \(t=m\) in \(f(s,t) = (3s-4t,2s+t)\) Thus, you should have that \(f(5m+n,m) = (3(5m+n) - 4(m), 2(5m+n)+(m))=\ldots\) Does this make sense?

OpenStudy (anonymous):

owh.. sorry.. im still weak in this topic.. i will do again as exercise.. I have one question that i still dont know how to answer.. here @ChristopherToni Let f: A → B be a function. Prove that f is bijective if and only if the inverse relation f^(-1) is a function from B to A. Give a numerical example.